What is the current economic value of an inheritance that will pay $3700 to the beneficiary at the beginning of every three months for 15 years starting when the beneficiary reaches 21 years of age, 5 1/4 years from now? Assume that money can earn 7.2% compounded monthly. (Do not round intermediate calculations and round your final answer to 2 decimal places.)

  Current economic value $

I need steps and not excel method
Answer #1

Firstly we will calculate the EAR

EAR = ((1+(Nominal rate/Number of compounding periods)^ Number of compounding periods)  - 1

EAR= [(1+(7.2%/12)) ^12] - 1

EAR = 7.44%

Now we will use a BA 2 Plus financial calculator to calculate the PV of $3700

I/Y (Interest per period) = 7.44%/ 4 = 1.86%

PMT (Payment per period) = $3700

N = 15*4 = 60

CMPT PV

PV = 135,563.4044

Current economic value of an inheritance that will pay $3700 to the beneficiary = 135,563.4044/(1.0744^5.25)

= $93,008.14